Grade 9Mechanics

A ball of 1 Kg strikes a wall with a velocity of 1m/s at an angle of 30degrees with the wall and reflects at the same angle.If it remains in contact with wall for 0.1 sec,then the force is

DSK
10 Years agoGrade 9

1 Answer

Anonymous
7 Years ago
Force= momentum ÷ time 
Total Momentum= m×u×cos30+m×v×cos30 (the angle with the normal)
=> P= 2*1*1*(√3/2)=√3
Force=√3÷0.1=10√3